XeF 4 + H 2 O=

XeF4 + H2O=



 

Grade:12

1 Answers

Ramesh V
14 years ago

XeF2(g) + H2O(l) = Xe(g) + HF(g) + O2(g)

balanced eqn is :   2XeF2(g) + 2H2O(l) = 2Xe(g) + 4HF(g) + O2(g)

--

regards

Ramesh